### 26 Dec 2023 v2

# Zhaojian Yu, Xin Zhang, Ning Shang, Yangyu Huang, Can Xu, Yishujie Zhao, Wenxiang Hu, Qiufeng Yin Microsoft

Abstract
Recent work demonstrates that, after being fine-tuned on a high-quality instruction dataset, the resulting model can obtain impressive capabilities to address a wide range of tasks. However, existing methods for instruction data generation often produce duplicate data and are not controllable enough on data quality. In this paper, we extend the generalization of instruction tuning by classifying the instruction data to 4 code-related tasks and propose a LLM-based Generator-Discriminator data process framework to generate diverse, high-quality instruction data from open source code. Hence, we introduce CodeOcean, a dataset comprising 20,000 instruction instances across 4 universal code-related tasks,which is aimed at augmenting the effectiveness of instruction tuning and improving the generalization ability of fine-tuned model. Subsequently, we present WaveCoder, a fine-tuned Code LLM with Widespread And Versatile Enhanced instruction tuning. This model is specifically designed for enhancing instruction tuning of Code Language Models (LLMs). Our experiments demonstrate that Wavecoder models outperform other open-source models in terms of generalization ability across different code-related tasks at the same level of fine-tuning scale. Moreover, Wavecoder exhibits high efficiency in previous code generation tasks. This paper thus offers a significant contribution to the field of instruction data generation and fine-tuning models, providing new insights and tools for enhancing performance in code-related tasks.
Figure 1: Comparision with CodeAlpaca with different dataset size(a) and different base models(b). CodeOcean outpeforms CodeAlpaca on HumanEval multidimensionally, more detailed analysis is shown in Section 4.3.
1 Introduction
Figure 2: The overview of the pipeline.
In recent months, numerous Large Language Models (LLMs) such as ChatGPT, GPT-4 [1], and Gemini  [^1] https://deepmind.google/technologies/gemini have attained unprecedented performance levels in a broad array of NLP tasks. These models utilize a self-supervised pre-training process, and subsequent fine-tuning, to demonstrate exceptional zero/few-shot capabilities, effectively following human instructions across an array of diverse tasks.
However, the overhead of holding such a large model is huge. Therefore, some relatively small LLM, especially Code Large Language Models (Code LLMs), have attracted considerable attention due to their exceptional performance on a wide range of code-related tasks. Given that LLMs can accrue a wealth of professional knowledge through pre-training [1, 2, 3], an efficient and effective pre-training process on a code corpus is critical for Code LLMs. Several previous works, including Codex [4], CodeGen [5], StarCoder [6], and CodeLLaMa [7], have successfully demonstrated that this process can significantly improve their capability to tackle code-related problems.
Additionally, instruction tuning has been shown to perform at a satisfactory level across various tasks, as demonstrated in research based on instruction-tuning like FLAN [8], ExT5 [9], and FLANT5 [10]. These studies incorporated thousands of tasks into their training pipelines to improve the generalization ability of pre-trained model on downstream tasks. InstructGPT [11] aligns user input effectively by integrating high-quality instruction data written by human labelers, which inspires further exploration into the utility of instruction tuning. Stanford Alpaca [12] utilized ChatGPT to generate instruction data via a self-instruct method, which in turn was used in the instruction tuning process. Further improvements to the effectiveness of pre-training models have been explored in WizardLM [13] and WizardCoder [14], where the evol-instruct method was applied. These works collectively exemplify the potential of instruction tuning in enhancing the performance of large language models.
Based on these work, our intuition is that instruction-tuning can activate the potential of LLMs then align the pre-trained models to an outstanding intelligent level. Therefore, we have summarized the main function of instruction tuning:
Generalization. Initially proposed to augment the cross-task generalization capabilities of Language Models (LMs)[8, 9, 10], instruction-tuning enhances performance across a multitude of unseen tasks when fine-tuned with diverse NLP task instructions.
Alignment. Pre-trained models [2, 3, 15, 16], having learned from a vast number of self-supervised tasks at both token and sentence levels, already possess the capability to comprehend text inputs. Instruction-tuning offers instruction-level tasks for these pre-trained models[15, 17], enabling them to extract more information from the instruction beyond the semantics of raw text. This additional information is user intention and enhances their interactive abilities with human users, thereby contributing to alignment.
To enhance the performance of Code LLMs through instruction tuning, many existing methods for generating instructional data have been devised, focusing primarily on the two aforementioned aspects. For example, self-instruct [12], evol-instruct [13] employed the zero/few-shot ability of teacher LLM to generated instruction data, which present an amazing way for instruction data generation. However, these generation ways are too dependent on the performance of the teacher LLM and sometimes produces a large amount of duplicate data which will reduce the efficiency of fine-tuning [18, 19, 20]. Octopack [21] constructs a code instruction dataset leveraging the natural structure of Git commits to break away from dependence on teacher LLMs. Howerver, Ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task.
To solve these problems, as shown in Figure 2, we propose a method that could make full use of source code and explicitly control the quality of generated data. Owing to the fact that instruction tuning is to align the pre-training model to the instruction-follow training set [1, 17], we prensent a LLM Generator-Disciminator framework for instruction data generation. By employing generation and discrimination, our method can make the data generation process more customizable and more controllable. Taking the raw code as input and selecting the core dataset, our method could stably generate more realistic instruction data and control the diversity of data by adjusting the distribution of the raw code.
In this paper, our focus is on enhancing the performance of code LLMs by applying instruction-tuning. Addressing the aforementioned challenges, we refine the instruction data by classifying the instruction instances to four universal code-related tasks: Code Summarization, Code Generation, Code Translation, and Code Repair and generate a dataset of 20,000 instruction instances, termed CodeOcean, for the four code-related tasks with our data generation strategy.
To validate our approach, taking StarCoder [6], CodeLLaMa [7], and DeepseekCoder [22] as our base models , we introduce WaveCoder models, fine-tuned with our data generation strategy and evaluate it on HumanEval [4], MBPP [23], HumanEvalPack [24] benchmarks, experimental results show that our WaveCoder exhibits exceptional performance based on a small-scale instruction tuning.
2 CodeOcean: Four-task Code-related Instruction Data
In this section, we introduce the methodological details of our exploration. We firstly choose 4 representative coding tasks and collect raw code from open source dataset. For each task, we employ GPT-3.5-turbo to generate instruction data for fine-tuning. The generation prompt is shown in Table 2 and the data generation process is shown in Figure 3.
Figure 3: The overview of the LLM-based Generator-Discriminator Framework.
2.1 Tasks Setting
Given the breadth of code-related tasks, we have selected four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration. These include Code Summarization, Code Generation, Code Translation, and Code Repair. Detailed descriptions of these tasks can be found below.
Table 1: The proportion of programming language in raw code.
Task Percentage(%)
Python 29.44
PHP 21.34
Go 19.68
Java 18.53
JavaScript 5.56
Others (Ruby,C++,C#) 5.45
Code Summarization (code-to-text). This task aims to create a brief summary of a given code. The raw code is used as input and the model’s response is reformulated into an instructional format.
Code Generation (text-to-code, code-to-code). In this task, the model is expected to generate code based on a user’s demand description. The model generates instructions to simulate the user’s demand discourse based on the raw code. The generated raw code is then considered as the output.
Code Translation (code-to-code). This task involves converting one programming language into another. A prompt and raw code are given to the model, and it generates instructions and the translated code.
Code Repair (code-to-code). The aim of this task is to provide correct code based on potential issues in the given code. The model is expected to generate solutions for the raw code, typically within one or two sentences, which are then taken as the output.
2.2 Generation of Training Data
In past research work[25, 26], many researchers have discovered that data quality and diversity play an important role in the instruction tuning process rather than data amount. The improvement of data quality and diversity are directly related to the performance of the fine-tuned model. Therefore, we introduce the details of our training data generation process in this section.
2.2.1 Selection of Raw Code
CodeSearchNet [^2] https://huggingface.co/datasets/code_search_net is a dataset containing 2 million (comment, code) pairs from open-source libraries hosted on GitHub. It includes code and documentation for 6 programming languages. We have chosen CodeSearchNet as our foundation dataset and applied the coreset-based selection method KCenterGreedy [27, 28] to maximize the diversity of the raw code.
In order to select high-quality code for instruction-tuning, we make some rules to filter the foundation dataset preliminarily. The main rules are as follows:
  • The length of required code should not be too long or too short. In this work, we filtered the code whose length lower than 50 or higher than 800.
  • Followed Code Alpaca [^3] https://github.com/sahil280114/codealpaca, We have eliminated the raw code containing words from the blacklist, since these code could potentially reduce the performance of the resulting model.
After filtered by the above rules, the size of the raw code dataset dropped to 1.2 million, which is still a huge scale and too expensive to fine-tune. To this end, we employed KCenterGreedy algorithm based on code embeddings to reduce the amount of training data while maximizing data diversity of raw code as much as possible.
In this method, our main work is divided into two steps:
Code embedding. In this step, we encode all the raw code samples using a pre-trained language model (roberta-large-v1[16] in this paper). we take the [CLS] token as the code embedding for one input raw code, which is not defined in generative language models like GPT-3.5.
Coreset sampling. In the coreset sampling step, our goal is to use a small number of samples to represent the distribution of the overall data, thus we employed unsupervised clustering and focused on the data representations. After obtaining the code embeddings from the first step, we used KCentergreedy algorithm [27] to select the core dataset and maximize data diversity, which aims to choose $k$ center points such that minimize the largest distance between a random data point and its nearest center, and has been proven efficient in obtaining a set of core samples of one distribution.
Table 2: The proportion of generated data and the prompts used.
Task Num Per(%) Prompt
Code Generation 11370 57.0 Implementing functions that perform specific operations given input.
Code Summarization 3175 15.9 Write clear and concise documentation for the given code.
Code Repair 3144 15.7 Identify and fix errors in the given code.
Code Translation 2236 11.2 Rewrite the given code from one programming language to another.
2.2.2 LLM-based Generator-Discriminator Framework
After the process of collecting raw code, it is significant to generate instruction data and reformat them to training data. To this end, shown in Figure 2, we propose a novel LLM-based Generator-Discriminator Framework that can leverage an extensive amount of unsupervised open source code to generate supervised instruction data. By incorporating the open source code, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself.
In the generation phase, we utilize GPT-4 to generate a task definition in the context of a specific scenario. We follow this with the definition of generation requirements for the task at hand, integrating both the task definition and all associated requirements into the generation prompt. By using the raw code as an input and selecting good and bad case examples, we leverage the few-shot abilities of GPT-3.5 to generate the necessary knowledge and information for instruction tuning.
For the discriminator phase, we establish a series of criteria related to instruction data and employ GPT-4 to assess the quality of instruction instances. Each instruction instance is classified as either a good or bad case, and this information is subsequently reused in the following generation as examples. An example of our Generator and Discriminator framework is presented in Figure 4. This framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality and diverse training dataset.
Figure 4: An example of LLM-based Generator and Discriminator framework. The generator produce the instruction data based on the input (a). Subsequently, the discriminator accept the output and generated analysis for it. The output (b) includes four keys, we take the information as input and solution as output in our instruction tuning. The analysis (c) consists of the detailed reason for each rule and an overall answer to check if the sample meet all the requirements.
3 Training
Unlike the previous work [14, 29, 30] that mainly focus on code generation task, we hope that CodeOcean could gain more general problem-solving capabilities on a wider range of coding tasks. Therefore, our training relies on a 20k instruction dataset generated by our LLM-based Generator-Discriminator framework and an additional logical dataset. The 20k dataset covers 4 common code tasks, including code generation, code translation, Code summarization, and code repair.
To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU. For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5. The sequence length of our each instruction sample is less than 1024. The fine-tuning prompt is as follows:
Prompt with Input: Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction:{instruction} ### Input:{input} ### Response: Prompt without Input: Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction:{instruction} ### Response:
4 Experiments
4.1 Experimental Setup
4.1.1 Benchmarks and Baselines
We evaluate our model on three code benchmarks: HumanEval [4], MBPP[23], HumanEvalPack[24].
HumanEval  [^4] https://huggingface.co/datasets/openai_humaneval, which consists of 164 manually-written Python programming problems and an average of 9.6 test cases allocated to each problem is now the most extensively adopted benchmark for Code LLMs.
MBPP  [^5] https://huggingface.co/datasets/mbpp consists of around 1,000 crowd-sourced Python programming problems, designed to be solvable by entry level programmers, covering programming fundamentals, standard library functionality, and so on. In this paper, we choose the 500 preoblems test dataset to evaluate both few-shot inference of fine-tuned models. For whose MBPP(500) result is not reported or not used, we reproduced for them using bigcode-evaluation-harness  [^6] https://github.com/bigcode-project/bigcode-evaluation-harness.
HumanEvalPack  [^7] https://huggingface.co/datasets/bigcode/humanevalpack is an extension of OpenAI’s HumanEval to cover 6 total languages across 3 tasks. In this paper, we select the HumanEvalFix to evaluate the code to code ability especially on code repair task and HumanEvalExplain benchmarks to evaluate the code to text ability especially on code summarization task.
Proprietary Models. We present the self-reported results from an array of SoTA LLMs, including Gemini, ChatGPT (gpt-3.5-turbo), GPT-4. If not reported, we use the results from Octopack or evaluate by ourselves.
Open source Models. Base models comprise StarCoder, CodeLLaMa and Deepseek-Coder with prompting. To ensure an equitable comparison, we opted to select those instructed models that have been trained using fewer than 100,000 instruction instances for our comparative analysis.
4.1.2 metrics
Pass@k. Followed most previous works, we also adopt the pass@k metric proposed by OpenAI [4] to evaluate the performance of our model. By generating $n$ samples $(n>k)$ for each promblem and checking the correctness for each sample with the test cases, the pass@k score can be calculated by the unbiased estimator:
$pass@k=E\left[1-\frac{{n\choose k-c}}{{n\choose k}}\right]$ (1)
where $n$ is the total number of sample generated for a problem, $c$ is the number of correct samples passing all the test cases.
4.2 Experimental Results
4.2.1 Evaluation on Code Generation Task
HumanEval and MBPP are two representative benchmarks for code generation task where the model should generate the complete code based on the function signature and the docstring of the problem. Table 3 shows the pass@1 score of different LLMs on both benchmarks. From the results, We have the following observations:
WaveCoder models substantially outperform the instructed models training with fewer than 20K instruction tuning data (InsT Data). Following the fine-tuning process, the performance of our models exhibit substantial improvement when compared to both the foundation model and a selection of open-source models, but it still lags behind proprietary models the instructed models training with more than 70K training data.
Refined and diverse instruction data can significantly improve the efficiency of instruction tuning. Although Wavecoder falls behind WizardCoder and Magicoder on code generation task, there is a considerable difference in data size among them. Nevertheless, the performance lag of Wavecoder is relatively minor.
Table 3: Results of pass@1 on HumanEval and MBPP benchmark. We use self-reported scores whenever available. Due to the difference in decoding strategies from previous evaluation work, we marked the results of greedy decoding in blue and n = 200 samples in red. -: Not reported in their paper.
Model Params Base Model InsT Data HumanEval MBPP (500)
Proprietary Models
GPT-4 [1] - - - 86.6 / 67.0 -
ChatGPT [1] - - - 73.2 / 48.1 52.2
Gemini - - - 74.4 -
Open-Source Models
StarCoder 15B - 33.6 43.3
OctoCoder [21] 15B StarCoder 13K 46.2 43.5
WizardCoder [14] 15B StarCoder 78K 57.3 51.8
WaveCoder-SC-15B [Uncaptioned image] 15B StarCoder 20K 50.5 (+16.9) 51.0 (+7.4)
CodeLLaMa [7] 7B - 33.5 41.4
CodeLLaMa-instruct [7] 7B CodeLLaMa 14K 34.8 44.4
WaveCoder-CL-7B [Uncaptioned image] 7B CodeLLaMa 20K 48.1 (+14.6) 47.2 (+5.8)
CodeLLaMa [7] 13B - 36.0 47.0
CodeLLaMa-instruct [7] 13B CodeLLaMa 14K 42.5 49.4
WaveCoder-CL-13B [Uncaptioned image] 13B CodeLLaMa 20K 55.4 (+19.4) 49.6 (+2.6)
DeepseekCoder 6.7B - 49.4 60.6
WaveCoder-DS-6.7B [Uncaptioned image] 6.7B DeepseekCoder 20K 62.8 (+13.4) 62.4 (+1.8)
Table 4: Results of pass@1 on HumanEvalFix benchmark. We use self-reported scores whenever available. Due to the difference in decoding strategies from previous evaluation work, we marked the results of greedy decoding in blue and n = 20 samples in red.
Model Python JavaScript Java Go C++ Rust Avg.
GPT-4 47.0 48.2 50.0 50.6 47.6 43.3 47.8
StarCoder 8.7 15.7 13.3 20.1 15.6 6.7 13.4
OctoCoder 30.4 28.4 30.6 30.2 26.1 16.5 27.0
WizardCoder 31.8 29.5 30.7 30.4 18.7 13.0 25.7
WaveCoder-SC-15B [Uncaptioned image] 39.3 35.1 34.8 36.2 30.2 22.5 33.0
CodeLLaMa-instruct-7B 28.0 23.2 23.2 18.3 0.1 0.1 15.5
CodeLLaMa-CodeAlpaca-7B 37.8 39.0 42.0 37.8 37.2 29.2 37.1
WaveCoder-CL-7B [Uncaptioned image] 41.4 41.4 42.0 47.1 42.7 34.7 41.5
CodeLLaMa-instruct-13B 29.2 19.5 32.3 24.4 12.8 0.1 19.7
CodeLLaMa-CodeAlpaca-13B 42.7 43.9 50.0 45.7 39.6 37.2 43.2
WaveCoder-CL-13B [Uncaptioned image] 48.8 48.2 50.6 51.8 45.1 40.2 47.4
DeepseekCoder-6.7B 29.9 29.2 39.0 29.2 25.0 21.9 29.0
DeepseekCoder-CodeAlpaca-6.7B 49.4 51.8 45.1 48.8 44.5 31.7 45.2
WaveCoder-DS-6.7B [Uncaptioned image] 54.9 51.2 57.3 49.3 45.1 32.9 48.4
4.2.2 Evaluation on Other Code-related Task
We score WaveCoder with state-of-the-art Code LLMs on HumanEvalPack in Table 4.We use the result reported in Octopack [21] when avaliable. Table 4 presents the results of WaveCoder on code repair task and Table 5 presents the results of WaveCoder on Code summarization task, highlighting the the following salient observations:
WaveCoder outperforms all open source models on other code-related task with similar degrees of fine-tuning. Using Starcoder as the foundation model, our WaveCoder-SC outperforms both WizardCoder and OctoCoder across all programming languages on HumanEvalFix and HumanEvalExplain benchmarks. Even when compared with other foundation models, our model consistently demonstrates superior performance over all previously fine-tuned models.
Enhancing the refinement and diversity of data can significantly improve the effectiveness of instruction tuning in multi-task scenarios. By refining the data and classifying the instructions to 4 different code-related downstream tasks, our models achieve a surprising performance level on both evaluation scenarios. Notably, WaveCoder-DS, utilizing only 6.7B parameters, approximates the performance of GPT-4 on HumanEvalFix quite closely.
Table 5: Results of pass@1 on HumanEvalExplain benchmark. We use self-reported scores whenever available. Due to the difference in decoding strategies from previous evaluation work, we marked the results of greedy decoding in blue and n = 20 samples in red.
Model Python JavaScript Java Go C++ Rust Avg.
GPT-4 64.6 57.3 51.2 58.5 38.4 42.7 52.1
StarCoder 0.0 0.0 0.0 0.0 0.0 0.0 0.0
WizardCoder 32.5 33.0 27.4 26.7 28.2 16.9 27.5
OctoCoder 35.1 24.5 27.3 21.1 24.1 14.8 24.5
WaveCoder-SC-15B [Uncaptioned image] 37.1 33.3 40.5 23.3 31.8 19.3 30.8
CodeLLaMa-instruct-7B 33.5 36.0 31.7 21.3 25.0 16.4 27.3
CodeLLaMa-CodeAlpaca-7B 34.7 24.4 37.8 23.2 28.6 19.5 28.0
WaveCoder-CL-7B [Uncaptioned image] 41.4 31.7 39.0 25.0 34.1 23.2 32.4
CodeLLaMa-instruct-13B 40.2 26.8 37.2 22.5 28.0 14.6 28.2
CodeLLaMa-CodeAlpaca-13B 32.3 28.0 34.1 18.9 29.9 20.7 27.3
WaveCoder-CL-13B [Uncaptioned image] 45.7 42.0 48.2 32.3 38.4 20.7 37.9
DeepseekCoder-6.7B 43.9 40.2 37.8 29.2 34.1 22.5 34.6
Deepseek-CodeAlpaca-6.7B 40.8 37.2 42.1 29.9 31.7 22.5 34.0
WaveCoder-DS-6.7B [Uncaptioned image] 47.5 45.7 48.2 33.5 41.4 27.4 40.6
4.3 Comparison with CodeAlpaca
CodeAlpaca dataset contains 20K multi-task instruction-following data generated by the techniques in the self-instruct [12]. To ensure a fair and multidimensional comparison, we randomly sampled 1K and 5K from both datasets (CodeAlpaca and CodeOcean), set the same set of training hyper-parameters set (epoch = 3, learning rate = 1e-4, LoRA rank = 8) and used the same training prompts. To prevent overfitting, we use Low-Rank Adaption (LoRA) [31] for fine-tuning if the size of instruction-follow training dataset is less than 5K and perform full fine-tuning on whole 20K dataset. Table 4, 5 present the result of CodeOcean on HumanEval and other code-related benchmarks highlighting the following salient observations:
  • After being fine-tuned with 1K, 5K and 20K of instructional data respectively, the performance of base model improves significantly on HumanEval shown in Figure 1.Taking Starcoder as the base model, CodeOcean surpasses the CodeAlpaca (44.9% vs 41.7%, 45.7% vs 48.1% and 47.0% vs 50.5%) shown in Figure 1 (a), which emphasizes the effectiveness of our LLM-based Generator-Discriminator Framework on refining instruction data. As shown in Figure 1 (b), The results of different base models on CodeOcean surpasses the the results on CodeAlpaca, which emphasizes the effectiveness of CodeOcean dataset in enhancing the instruction-following ability of the base model.
  • According to Table 4 and Table 5, All Wavecoder models significantly outperform the model fine-tuned with CodeAlpaca. Remarkably, The pass@1 score of Wavecoder-CL-13B outperforms CodeLLaMa-CodeAlpaca-13B achieving 10.6% absolute improvements on HumanEvalExplain. This emphasizes the effectiveness of defining and classifying code-related tasks on enhancing the generalization ability of Code LLMs.
5 Related Work
5.1 Code Large Language Models.
Numerous researchers have proposed their Code Language Models (LLMs), including CodeGen [32], CodeT5[33], StarCoder [6], CodeLLaMa [7], and Deepseek-Coder [22] within coding domain. By leveraging a comprehensive pre-training process with a vast amount of code knowledge from scratch, these Code LLMs often exceed the performance of super LLMs in many scenarios. In a bid to generate output more efficiently and assist in resolving programming issues, instruction-based models like InstructCodeT5+ [34] and WizardCoder [14] have been introduced.
5.2 Instruction Tuning.
In order to unleash the potential of the foundation LLMs, instruction tuning is proposed to fine-tune the LLM with multi-task datasets which is presented in task descriptions by natural language. For example, OpenAI fine-tuned its GPT-3 [3] model with a diverse instruction dataset written by human annotators, the instruction-tuned model can effectively align the intention of human users. This work is the indispensable part of the InstructGPT [17], which is the foundation of the well-known ChatGPT. [10] proposes Flan-PaLM 540B [10] which is instruction-tuned on 1.8K tasks and finds that the diversity of instruction dataset can significantly improve the performance of the instruction-tuned language model. Alpaca [12] adopts the self-instruct method, which leverages the few shot ability of ChatGPT to generate data for training. For Code LLMs, InstructCodeT5+ [34], WizardCoder [14], Pangu-coder2 [29], are recent recent models trained with coding instructions. However, all the instruction data they used is from CodeAlpaca, which is not refined enough.
6 Conclusion
This paper studies the effect of multi-task instruction data on enhancing the generalization ability of Code Large Language Model. We introduce CodeOcean, a dataset with 20K instruction instances on four universal code-related tasks. This method and dataset enable WaveCoder, which significantly imporove the generalization ability of foundation model on diverse downstream tasks. Our WaveCoder has shown the best generalization ability among other open-source models in code repair and code summarization tasks, and can maintain high efficiency on previous code generation benchmarks. In order to further improve the mono-task performance and generalization ability of the model, future work should focus on the interplay among different tasks and larger dataset.